Wendy was awarded a volleyball scholarship to the University of Michigan, so on graduation her parents gave her the $15,000 they had saved for her college tuition. She opted to invest some money in a privately held company that pays 10% per year and evenly split the remaining money between a market account yielding 2% and a high rick stock that yielded 40%. At the end of the first year she had $17,600 total. How much did she invest in each of the three?
10%=
2%=
40%=

Answer by Boreal(15235)

She earned $2600
x was invested at 10% earning .10x
the rest was 15000-x
split means (1/2)(15000-x)
between 2% and 40%
so .02(.5)(15000-x)=.01(15000-x)
and (.40)(.5)(15000-x)=0.2(15000-x)
2600=.10x+.150-.01x+3000-.2x
2600=-.11x+3150
-550=-.11x
x=$5000 at 10% earning $500
$5000 at 2% earning $100
$5000 (the other half of the split) earning 40% or $2000.
That is $2600.